The area of a triangle is one hald the base times the height. If the height of a triangle is 4cm and the area is 40 sq. cm, then what is one half the length of the base?
Please also include the formula you used.
thank you

Answer by fcabanski(1391)

A = (1/2 * L) * h


You want to find 1/2 * L, so divide both sides by h.


A/h = 1/2 * L


40/4 = 10cm = 1/2 * L.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)